During a biological process, 3900 K+ ions pass through a channel in the phospholipid bilayer of a cell membrane in 680 μs.

a) What is the electric current through the channel during that time interval?

b) If the potential difference across the channel is 75 mV, what electric power is associated with the current?
1
Expert's answer
2020-07-03T15:01:55-0400

Number of the ions passing through the channel = 3900

time of the flow of the ions = 680μs

(i) Current through the channel, "I = \frac{Charge }{Time} = \frac{3900*1.6*10^{-19}}{680*10^{-6}} = 9.17*10^{-13} A"

(ii) Potential Difference applied = 75 mV

Power associated, "P = VI = 9.17*10^{-13} *75*10^{-3} = 687.75 * 10^{-16} J"